OpenStudy (anonymous):

Write the equation of the circle with center at (0 , 0) and a radius of 6.

OpenStudy (saifoo.khan):

\[(x - 0)^2 + (y - 0)^2 = 6^2\]Solve.

OpenStudy (anonymous):

Sorry i dont need the answer i want is the formula or the process

OpenStudy (anonymous):

\[(x - h)^2 + (y - k)^2 = r^2\] Where, (h,k) = Centre and r = Radius.. \[(x-0)^2 + (y - 0)^2 = 36\] And solve...

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

\[x ^{2}+y ^{2}= 36\]

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

is this the answer?

OpenStudy (anonymous):

your right. good.